# zkFuzz: Foundation and Framework for Effective Fuzzing of Zero-Knowledge Circuits

### v1

@@ 6.1. RQ1: Effectiveness of zkFuzz
Tab. 4 summarizes the total number of unique bugs detected by each tool across different benchmark categories. We manually check all potential bugs flagged by at least one tool and categorize them into TP (True Positive), which refers to cases where the tool correctly flags ZK circuits as unsafe, and FP (False Positive), which refers to cases where the tool incorrectly flags a test suite as unsafe. Precision is calculated as TP / (TP + FP). Since there may be undetected bugs not flagged by any tool, we do not report true or false negatives. The #Bugs column shows the number of unique vulnerabilities detected by at least one tool. A bug in a Circom template (§ 2.3) may affect all programs that instantiate the template, but we count it as a single bug to avoid double counting. Moreover, when a vulnerability is detected in multiple categories (for example, if a bug is identified in both a small and a medium circuit because the use the same buggy template), it is counted only in the smallest applicable category to prevent double counting. Note this way of counting favors existing formal tools due to their limited scalability. We report the maximum and minimum TP of zkFuzz with five different random seeds. Our evaluation shows that zkFuzz significantly and stably outperforms existing methods, detecting about 96% (66/69) of all bugs without any false positives. While very large templates present a bottleneck for dynamic testing, causing zkFuzz to underperform Circomspect in the ”Very Large” category slightly, zkFuzz still identifies 66% (4/6) of the bugs.
@@
In addition, zkFuzz-Naive, which performs the search without any guidance, still detects more bugs than other existing tools, demonstrating the strength of our core approach. The default zkFuzz detects 10% more bugs overall compared to zkFuzz-Naive, especially in larger circuits.
@@
Our benchmark shows that the performance of ZKAP is worse than that reported in (Wen et al., 2024). This stems from not classifying potential logic issues, such as unused inputs, as vulnerabilities. Furthermore, the official implementation of ZKAP is known to crash on certain templates written in recent Circom versions—an issue acknowledged by ZKAP’s developers but yet to be fixed. The results of ZKAP with US and USCO detectors and our modified version of Circomspect with recursive analysis are provided in Appendix D.
@@
Our analysis also reveals that 53.5% of root causes stem from computation abort bugs, while 46.5% are due to non-deterministic behavior. This result highlights the importance of comprehensively addressing both types of bugs, whereas Picus and ConsCS focus solely on non-deterministic cases.
@@
ConsCS detects fewer bugs overall compared to Picus. Although Jiang et al. (Jiang et al., 2025) reports a higher success rate for ConsCS, it is important to note that their evaluation measures both safe and unsafe circuits (demonstrating either security or vulnerability). In contrast, our experiment focuses solely on vulnerability detection and does not include formal verification of circuit safety.
@@
Result 1: Effectiveness of zkFuzz zkFuzz has efficiently detected 66 out of 69 vulnerabilities in the benchmark of real-world ZK circuits without any false positives, demonstrating its reliable ability to identify critical security flaws in ZK circuits.
6.2. RQ2: Detection Speed
@@
Figure 3. Detection Time Analysis. For zkFuzz, we plot the average and standard deviation from five different random seeds. zkFuzz is superior to other methods.
@@
Fig. 3 illustrates the cumulative number of unique vulnerabilities detected by different tools over time. For zkFuzz, we depict the average and 1-sigma regions across five trials with different random seeds. Notably, zkFuzz identifies over 90% of bugs within the first 100 seconds, although zkFuzz-Naive takes more than one hour to achieve the same number of detected bugs, showing the importance of guided search with our fitness function and target selectors. Although ZKAP is a static analyzer, it operates more slowly than zkFuzz due to the computational overhead of compiling Circom files to LLVM and analyzing its dependency graph. Similarly, Picus exhibits slower performance, as it relies on expensive queries to SMT solvers. Our results show that ConsCS finds bugs faster than Picus confirming the findings of Jiang e al. (Jiang et al., 2025). The dashed line of Circomspect indicates that it finishes the analysis of all test cases before the timeout.
@@
Result 2: Detection Speed of zkFuzz zkFuzz identifies 90% of bugs within just 100 seconds, demonstrating its practical effectiveness over traditional formal methods. The comparison with zkFuzz-Naive highlights the advantage of guided search, achieving a speedup of more than 30 times.
6.3. RQ3: Previously Unknown Bugs
@@
Out of 66 identified bugs zkFuzz, 38 were previously unknown. Of these, 18 were found in production-level services and deployed smart contracts, 18 were confirmed by developers, and 6 have already been fixed. One example is shown in Code 8, found in the passport-zk-circuits (Rarimo, 2025), a Web3 project that raised \$2.5 million. For instance, when day=4, the expected assignment is {dayD:0, dayR:4} although {dayD:0, dayR:-6} also satisfies the constraint.
@@
7 dayD * 10 + dayR $=$ day;
8 signal dayE <== (dayD * 2**8 + dayR) + (2**4 + 2**5 + 2**12 + 2**13);
9//Encode month and year in the same manner
@@
Code 8: A previously unknown bug detected by zkFuzz in passport-zk-circuits (Rarimo, 2025). This circuit is under-constrained.
Our zkFuzz can generate counterexamples, unlike static analyzers such as ZKAP and Circomspect, significantly reducing the need for manual inspection after automated vulnerability detection. Among the six fixed vulnerabilities identified by zkFuzz, the average time to merge a fix after reporting was just 4.3 days—10 to 20 times faster than the typical resolution time for accepted security issues in open-source projects (Bühlmann and Ghafari, 2022; Rebatchi et al., 2024). As of now, no developer has concluded that a report from zkFuzz was a false positive.
Result 3: Previously Unknown Bugs found by zkFuzz zkFuzz can uncover 38 previously unknown bugs in production-grade ZK circuits, with 18 of them found in production-level services and deployed smart contracts.
6.4. RQ4: Ablation Study of Target Selectors
@@
Tab. 5 shows the results of ablation studies where we assess the impact of removing individual target selectors: error-based fitness score, skewed distribution, white list, invalid array subscript, hash-check and zero-division. We report the average and standard deviation of the cumulative number of unique bugs found at each time checkpoint among five different random seeds. For skewed distribution, we use a less skewed distribution; binary values (0 and 1) with 15%, small positive integers (2 to 100) with 34%, larger values near $q$ ($q-1000$ to $q-1$) with 50%, and all other values (11 to $q-1001$) with 1%. Fig. 4 also presents the relative execution time for discovering each unique bug, normalized against the default configuration.
@@
Among these factors, the choice of skewed distribution has the largest initial impact. Our default setting (Ours) finds 10 more bugs within the first 100 seconds compared to the less skewed distribution. The relative execution time indicates that adopting a less skewed distribution or removing the whitelist slows down detection by over 100×. In contrast, removing checks for invalid array subscripts, hash integrity, or division by zero prevents the fuzzer from detecting several bugs. However, after a reasonable amount of time, the performance of the less skewed distribution also converges to a similar number of unique detected bugs, demonstrating the stability and robustness of zkFuzz regardless of the specific choice of skewed distributions. Additionally, removing static analyzers, such as checks for invalid array subscripts, hash mismatches, and division by zero, results in several bugs being missed, highlighting their importance in the system’s overall effectiveness.
@@
Table 5. Impact of removing individual target selector on zkFuzz’s performance. Each selector contributes significantly to the number of detected bugs and the detection speed, demonstrating their importance.
Time (s) Ours
Less Skewed
Distribution
w.o. White List
w.o. Invalid Array
Subscript
w.o. Hash Check w.o. Zero Division
0.1 23.40 ± 1.02 18.20 ± 1.94 22.40 ± 2.87 25.00 ± 3.03 24.00 ± 1.67 20.80 ± 2.71
1.0 42.60 ± 4.50 35.20 ± 0.40 43.00 ± 2.53 43.40 ± 5.04 42.00 ± 5.06 41.40 ± 4.45
10.0 51.40 ± 3.44 43.00 ± 1.41 47.60 ± 2.33 49.60 ± 4.08 49.40 ± 3.83 50.40 ± 3.56
100.0 61.80 ± 0.75 49.20 ± 0.75 57.80 ± 2.23 60.00 ± 0.89 59.80 ± 0.75 60.80 ± 0.75
1000.0 63.60 ± 0.49 54.80 ± 2.48 62.20 ± 0.75 62.00 ± 0.63 61.60 ± 0.49 62.60 ± 0.49
3600.0 65.00 ± 0.63 62.00 ± 2.10 64.40 ± 0.49 63.20 ± 0.75 63.00 ± 0.00 64.00 ± 0.63
7200.0 65.40 ± 0.49 64.00 ± 0.89 64.60 ± 0.49 63.40 ± 0.49 63.00 ± 0.00 64.60 ± 0.49
@@
Figure 4. Relative execution time for discovering each unique bug compared to the default setting. Removing each heuristic degrades detection performance by over 100x.
@@
Result 4: Individual Contribution of Each Target Selector Each target selector in zkFuzz significantly enhances bug detection by increasing detection speed by more than 100 times and enabling the discovery of bugs that cannot be found without them.
6.5. RQ5: Hyperparameter Sensitivity
We further examine how mutation strength and population size affect performance. Fig. 5 compares configurations using weak mutation (mutation and crossover probabilities of 0.1) versus strong mutation (both set to 0.7) and small population (the number of program mutants and generated inputs of 10) versus large population (both set to 50). Our results indicate that weak mutation and a large population yield slightly inferior performance compared to the default setting, showing the importance of sufficient exploration in navigating the vast search space since weak mutation slows the speed of exploration while larger populations cause the fuzzer to execute more programs before steering mutants toward promising directions, increasing overhead. Nonetheless, regardless of hyperparameter choices, cumulative detections converge to similar levels.
@@ 6.5. RQ5: Hyperparameter Sensitivity
Figure 5. Impact of mutation strength and population size on zkFuzz. zkFuzz consistently maintains strong bug detection performance across various hyperparameter settings.
@@
Result 5: Stability of zkFuzz zkFuzz’s bug detection performance remains stable across different hyperparameter choices.
Ethical Considerations. During our research, we identified several previously unknown vulnerabilities in public projects. To ensure responsible disclosure, we contacted the project developers directly via email or social media whenever possible. We opened GitHub issues if we received no response or direct contact wasn’t available. All vulnerability reports were submitted at least 30 days before this paper, with most sent over 60 days prior.
7. Related Work
@@
Zero-knowledge proofs are vital for privacy and trust in decentralized systems, including cryptocurrencies and smart contracts. However, flaws in circuit constraints can compromise security by permitting malicious proofs or rejecting valid ones (Tang et al., 2024; Liang et al., 2025). We summarize different approaches to automatically detect various types of ZKP bugs below.
@@
Static Analysis. These approaches are widely used but suffer from high false positives. For example, Circomspect (of Bits, 2024) flags weak assignments (<--) indiscriminately, leading to many false positives. ZKAP (Wen et al., 2024) improves accuracy using circuit dependence graphs, yet it still identifies unconstrained inputs as potential vulnerabilities, which may be expected behavior. Despite identifying issues, static analyzers can’t generate counterexamples, requiring manual verification of flagged bugs.
@@
Formal Methods. Technique using formal methods rigorously verify ZK circuit correctness but struggle with scalability and require manual annotations. Tools like CIVER (Isabel et al., 2024) and Constraint Checker (Fan et al., 2024) need human input, while others like CODA (Liu et al., 2024) require rewriting circuits in another language. Fully automatic tools like Picus (Pailoor et al., 2023) and ConsCS (Jiang et al., 2025) use SMT solvers but still face scalability challenges and offer narrower definitions of under- and over-constrained circuits than ours. For example, they don’t capture under-constrained circuits caused by unexpected inputs (Wen et al., 2024).

### v2

@@ 6.1 RQ1: Effectiveness of zkFuzz
Tab. IV summarizes the total number of unique bugs detected by each tool across different benchmark categories. We manually check all potential bugs flagged by at least one tool and categorize them into TP (True Positive), which refers to cases where the tool correctly flags ZK circuits as unsafe, and FP (False Positive), which refers to cases where the tool incorrectly flags a test suite as unsafe. Precision is calculated as TP / (TP + FP). The lower gray row in each category corresponds to the results on the ZKAP dataset only [13]. Since there may be undetected bugs not flagged by any tool, we do not report true or false negatives. The #Bugs column shows the number of unique bugs detected by at least one tool. A bug in a Circom template (§ 2.3) may affect all programs that instantiate the template, but we count it as a single bug to avoid double-counting. Moreover, when a vulnerability is detected in multiple categories (for example, if a bug is identified in both a small and a medium circuit because they use the same buggy template), it is counted only in the smallest applicable category to prevent double-counting. Note this way of counting favors existing formal tools due to their limited scalability. We report the maximum and minimum TP of zkFuzz with five different random seeds. Our evaluation shows that zkFuzz++ significantly and stably outperforms existing methods, detecting about 96% (85/88) of all bugs without any false positives.
@@
In addition, zkFuzz, which performs the search without static analysis, still detects more bugs than other existing tools, demonstrating the strength of our core approach. zkFuzz++ detects 10% more bugs overall compared to zkFuzz, especially in larger circuits.
One clear advantage of zkFuzz over static analyzers like Circomspect and ZKAP is its ability to avoid false positives. This makes zkFuzz more practical as a testing tool, as developers are often inclined to disregard alerts from tools with high false positive rates [55, 56].
@@
Our benchmark shows that the performance of ZKAP is worse than that reported in [13]. This stems from not classifying potential logic issues, such as unused inputs, as vulnerabilities. Furthermore, the official implementation of ZKAP is known to crash on certain templates written in recent Circom versions—an issue acknowledged by ZKAP’s developers but yet to be fixed. The results of ZKAP with US and USCO detectors and our modified version of Circomspect with recursive analysis are provided in Appendix D.
@@
Our analysis also reveals that 55.2% of root causes stem from computation abort bugs, while 44.8% are due to non-deterministic behavior. This result highlights the importance of comprehensively addressing both types of bugs, whereas Picus and ConsCS focus solely on non-deterministic cases.
@@
ConsCS detects fewer bugs overall compared to Picus. Although [25] reports a higher success rate for ConsCS, it is important to note that their evaluation measures both safe and unsafe circuits (demonstrating either security or vulnerability), while we focus solely on vulnerability detection and do not include formal verification of circuit safety.
@@
Both zkFuzz and zkFuzz++ outperform existing tools, efficiently detecting 85 of $8$ vulnerabilities in the benchmark of real-world ZK circuits without false positives, demonstrating its reliable ability to identify critical security flaws in ZK circuits.
6.2 RQ2: Detection Speed
@@
Figure 3: Detection Time Analysis. Both modes of zkFuzz are superior to other methods.
@@
Fig. 3 shows the cumulative number of unique vulnerabilities detected by each tool over time. For zkFuzz and zkFuzz++, we depict the average and 1-sigma regions across five trials with different random seeds. Notably, zkFuzz++ identifies over 90% of bugs within the first 100 seconds, although the naive zkFuzz takes more than one hour to achieve the same number of detected bugs, showing the effectiveness of our target selectors. Although ZKAP is a static analyzer, it suffers from higher performance overhead than zkFuzz due to the overhead of compiling Circom files to LLVM and analyzing its dependency graph. Similarly, Picus exhibits slower performance, relying on expensive queries to SMT solvers. Our results show that ConsCS finds bugs faster than Picus, confirming the findings of [25]. The dashed line of Circomspect shows that it finishes the analysis of all test cases before the timeout.
@@
zkFuzz++ identifies 90% of bugs within just 100 seconds, demonstrating its practical effectiveness over traditional formal methods. The comparison with zkFuzz highlights the advantage of target selectors, achieving a speedup of more than 30 times.
6.3 RQ3: Previously Unknown Bugs
@@
Out of 85 bugs identified by zkFuzz++, 59 were previously unknown. Of these, 39 were confirmed by developers, and 14 have already been fixed. One example is shown in Code 8, found in the passport-zk-circuits [32], a Web3 project that raised \$2.5 million. For instance, when day=4, the expected assignment is {dayD:0, dayR:4} although {dayD:1, dayR:-6} also satisfies the constraint.
@@
6 dayD * 10 + dayR === day;
7 signal dayE <== (dayD * 2**8 + dayR) + 12336;
8//Encode month and year in the same manner
@@
Code 8: A previously unknown bug detected by zkFuzz in passport-zk-circuits [32]. This circuit is under-constrained.
Another example in Code 9 detected in the zk_whitelist [57]. This circuit is expected to abnormally terminate when the two inputs are different by assert, while this condition is not included in the constraints. This bug, confirmed by the developer and fixed, does not align with any prior formulations or patterns of ZK circuit vulnerabilities, showing the superiority of our TCCT. The same misuse of assert is also found in two circuits of masa-zkSBT, confirmed the developer.
1template Whitelist() {
2 signal input addressInDecimal, publicAddress;
3 signal output c;
4 component ise = IsEqual();
5 ise.in[0] <== addressInDecimal;
6 ise.in[1] <== publicAddress;
7 assert(ise.out==1);
8 c <== ise.out; }
Code 9: A previously unknown bug found by zkFuzz in zk_whitelist [57]. Any existing tools cannot capture this bug.
Unlike static analyzers such as ZKAP and Circomspect, zkFuzz can generate concrete counterexamples, significantly reducing the need for manual inspection. The bugs identified by zkFuzz were patched within a median of just 5.0 days after disclosure, underscoring both their critical severity and the prompt response of the maintainers.
zkFuzz++ uncovers 59 previously unknown bugs in real-world circuits, including ones missed by all existing tools due to incomplete formulations.
6.4 RQ4: Ablation Study of Target Selectors
@@
Tab. V shows the results of ablation studies where we assess the impact of removing individual target selectors: error-based fitness score, skewed distribution, white list, invalid array subscript, hash-check and zero-division. We report the average and standard deviation of the cumulative number of unique bugs found at each time checkpoint among five different random seeds. For skewed distribution, we use a less skewed distribution; binary values (0 and 1) with 15%, small positive integers (2 to 100) with 34%, larger values near $q$ (q1000q-1000 to q1q-1) with 50%, and all other values (11 to q1001q-1001) with 1%. Fig. 4 also presents the relative execution time for discovering each unique bug, normalized against the default configuration.
@@
Among these factors, the choice of skewed distribution has the largest initial impact. Our default zkFuzz++ (Ours) finds 10 more bugs within the first 100 seconds compared to the less skewed distribution. The relative execution time indicates that adopting a less skewed distribution or removing the whitelist slows down detection by over 100×. In contrast, removing checks for invalid array subscripts, hash integrity, or division by zero prevents the fuzzer from detecting several bugs. However, after a reasonable amount of time, the performance of the less skewed distribution also converges to a similar number of unique detected bugs, demonstrating the stability and robustness of zkFuzz regardless of the specific choice of skewed distributions. Additionally, removing static analysis-based target selectors results in several bugs being missed, highlighting their importance in the system’s overall effectiveness.
@@
TABLE V: Impact of removing individual target selector on zkFuzz++’s performance. Each selector contributes significantly to the number of detected bugs and the detection speed, demonstrating their importance.
Time (s)
Ours
(zkFuzz++)
Less Skewed Distribution w.o. White List w.o. Invalid Array Subscript w.o. Hash Check w.o. Zero Division
10.0 71.40 ± 3.44 61.60 ± 1.62 66.40 ± 3.01 69.60 ± 4.08 68.40 ± 3.83 70.40 ± 3.56
100.0 81.80 ± 0.75 68.80 ± 0.98 77.00 ± 2.19 80.00 ± 0.89 78.80 ± 0.75 80.80 ± 0.75
1000.0 83.60 ± 0.49 74.80 ± 2.48 82.20 ± 0.75 82.00 ± 0.63 80.60 ± 0.49 82.60 ± 0.49
7200.0 85.40 ± 0.49 84.00 ± 0.89 84.60 ± 0.49 83.40 ± 0.49 82.00 ± 0.00 84.60 ± 0.49
@@
Figure 4: Relative execution time for discovering each unique bug compared to the default setting. Removing each heuristic degrades detection performance by over 100x.
@@
Each target selector in zkFuzz++ significantly enhances bug detection by increasing detection speed by more than 100 times and enabling the discovery of bugs that cannot be found without them.
6.5 RQ5: Hyperparameter Sensitivity
@@
Figure 5: Impact of mutation strength and population size on zkFuzz++. zkFuzz++ consistently maintains strong bug detection performance across various hyperparameter settings.
We further examine how mutation strength and population size affect performance. Fig. 5 compares configurations of zkFuzz++ using weak mutation (mutation and crossover probabilities of 0.1) versus strong mutation (both set to 0.7) and small population (the number of program mutants and generated inputs of 10) versus large population (both set to 50). Our results show that weak mutation and a large population yield slightly inferior performance compared to the default setting, showing the importance of sufficient exploration in navigating the vast search space since weak mutation slows the speed of exploration while larger populations cause the fuzzer to execute more programs before steering mutants toward promising directions, increasing overhead. Nonetheless, regardless of hyperparameter choices, cumulative detections converge to similar levels.
@@
The bug detection performance of our fuzzer remains stable across different hyperparameter choices, identifying the same number of bugs in all settings.
6.6 RQ6: Preliminary Prototype for Noir
To demonstrate the language-agnostic capabilities of zkFuzz, we target Noir as a representative modern ZK DSL, which is a next-generation zero-knowledge (ZK) DSL, gaining traction. While Noir can automatically generate constraints for witness computation, developers often write manual constraints to optimize circuit size—enabling more efficient ZK circuits but also introducing opportunities for subtle bugs.
Supporting Noir in zkFuzz presents substantially more difficult challneges than Circom, as Noir’s rich, Rust-inspired syntax—with its advanced type system and user-defined abstractions—demands much more sophisticated parsing and emulation infrastructure, all within an ecosystem that is still in its early stages of development.
Therefore, our preliminary prototype of zkFuzz for Noir leverages its intermediate representations (IRs). Specifically, the Noir compiler lowers the Rust‐like program into two IRs: ACIR, which corresponds to the arithmetic circuit constraints, and Brilling, which corresponds to the unconstrained witness computation logic. As our mutation strategy, we modify the source operand of the Mov instructions in the Brilling opcodes to reference special memory regions; our custom memory structure then returns random values with the appropriate data type when those regions are read. In this way, we can mutate the witness‐calculation program without altering the underlying constraints. Although we have not yet implemented any target selectors for this Noir prototype, doing so is clearly feasible with a modest amount of engineering effort—an avenue we leave to future work.
Since there is no existing fuzzers for Noir targeting under or over-constrained circuit or dataset of vulnerable Noir circuits, we evaluate the preliminary implementation of zkFuzz for Noir on 20 representative Noir circuits from 8 open-source libraries, chosen by ourselves based on the number of stars and forks on GitHub and compatibility with the latest version of Noir compiler, where zkFuzz successfully rediscovers a known under-constrained bug [58].
Our preliminary prototype for Noir (which no other tools support) successfully identified an under-constrained circuit, showing that zkFuzz’s potential to generalize beyond a single language.
Ethical Considerations. We responsibly disclosed all previously unknown vulnerabilities found in public projects. When possible, we contacted maintainers privately; otherwise, we opened GitHub issues. Most disclosures were made at least 60 days before submission, and the detailed information about any unfixed bugs discovered after that cutoff is not included in this submission.
7 Related Work
@@
Zero-knowledge proofs are vital for privacy and trust in decentralized systems, including cryptocurrencies and smart contracts. However, flaws in circuit constraints can compromise security by permitting malicious proofs or rejecting valid ones [17, 15]. We summarize different approaches to automatically detect various types of ZKP bugs below (see Appendix E for the comparison table).
@@
Static Analysis. These approaches are widely used but suffer from high false positives. For example, Circomspect [20] flags weak assignments (<--) indiscriminately, leading to many false positives. ZKAP [13] improves accuracy using circuit dependence graphs, yet it still identifies unconstrained inputs as potential vulnerabilities, which may be expected behavior. Despite identifying issues, static analyzers can’t generate counterexamples, requiring manual verification of flagged bugs.
@@
Formal Methods. Formal methods rigorously verify ZK circuit correctness but struggle with scalability and require manual annotations. Tools like CIVER [22] and Constraint Checker [29] need human input, while others like CODA [23] require rewriting circuits in another language. Fully automatic tools like Picus [18] and ConsCS [25] use SMT solvers but still face scalability challenges and offer narrower definitions of under- and over-constrained circuits than ours. For example, they don’t capture under-constrained circuits caused by unexpected inputs [13].